OpenStudy (anonymous):

( 2 + √2 +1 / 2+√2+ 1 / 2 -√2 ) simplifies to:- a) 2-√2 b) 2 c)2+√2 d) 2√2 how, whats the method?

OpenStudy (anonymous):

@ shubhamsrg..plse help

OpenStudy (anonymous):

@shubhamsrg

OpenStudy (ajprincess):

Is the question \[2+\sqrt2+\frac{1}{2}+\sqrt2+\frac{1}{2}-\sqrt2\]?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

2+2/2+√2 =2+1+√2 =3+√2

OpenStudy (anonymous):

is that answer

OpenStudy (ajprincess):

ya that's the answer if what i have written is ur question:)

OpenStudy (anonymous):

but the answer options are a) 2-√2 b) 2 c)2+√2 d) 2√2

OpenStudy (ajprincess):

R u sure u didnt miss any - sign in front of one of the halves?

OpenStudy (anonymous):

yes ..i m sure